Find the p-value of an f-statistic using excel?

Hello, I am trying to understand how to compute the p-value of an f-statistic using excel using the data below I have found the f-statistic to be : 45.9210236


Jensen Tire & Auto is in the process of deciding whether to purchase a maintenance contract for its new computer wheel alignment and balancing machine. Managers feel that maintenance expense should be related to usage, and they collected the following information on weekly usage (hours) and annual maintenance expense (in hundreds of dollars).How to

Weekly Usage
(hours)
Annual
Maintenance Expense
18 20
15 25
25 33
33 40
37 50
22 34
29 36
36 42
45 55
43 43
  1. Develop the estimated regression equation that relates annual maintenance expense (in hundreds of dollars) to weekly usage hours (to 3 decimals).
    Expense =  +  Weekly Usage
  2. Test the significance of the relationship in part (a) at a .05 level of significance.

    Compute the value of the F test statistic (to 2 decimals).
Answer #1

here degree of freedom for numerator =number of independent variable =1

degree of freedom for denominator =n-2=10-2 =8

now for test statistic 45.92 and (1,8) degree of freedom , from excel Fdist function":

p value =Fdist(45.92,1,8) =0.0001

(please reply if help is required in any other part of this)